OpenStudy (anonymous):

Indicate what variations occur by marking as follows: A: Direct variation B: Inverse variation C: Joint variation D: Combined Variation E: No variation. 1. y=3x 2. xy=14 3. x^2=9 4. y=6xz 5. a=b/5 6. r=2s/t 7. x-1=10 8. 36=x/y Can you explain the terms above to me so i can try and do these 8 question Please.

jimthompson5910 (jim_thompson5910):

Direct variation: as the independent variable goes up, the dependent variable goes up. Likewise, this applies in the opposite direction: as the independent variable goes down, the dependent variable goes down.

jimthompson5910 (jim_thompson5910):

So this means that direct variation is in the basic form y = kx for some constant k

jimthompson5910 (jim_thompson5910):

So which equations do you see in this form?

OpenStudy (anonymous):

Number one.

jimthompson5910 (jim_thompson5910):

exactly

jimthompson5910 (jim_thompson5910):

any others?

OpenStudy (anonymous):

Number 4?

jimthompson5910 (jim_thompson5910):

no, it's close though...the only problem is that there are 2 variables on the right side instead of 1

OpenStudy (anonymous):

Oh, ok. I see that now.

jimthompson5910 (jim_thompson5910):

can we rewrite any equations and get them into a more standard form?

OpenStudy (anonymous):

Um, im not sure how to.

jimthompson5910 (jim_thompson5910):

that matches y = kx

jimthompson5910 (jim_thompson5910):

alright, notice that in #5, we're dividing by a scalar

jimthompson5910 (jim_thompson5910):

since multiplication is the opposite of division, this means that we can represent b/5 as 1/5 times b

jimthompson5910 (jim_thompson5910):

so in #5, a = b/5 turns into a = 1/5 times b

jimthompson5910 (jim_thompson5910):

does that fit the form y = kx?

OpenStudy (anonymous):

Oh ok. So that would make it a Direct variation?

jimthompson5910 (jim_thompson5910):

yes you are correct

OpenStudy (anonymous):

I dont think there is anymore, is there?

jimthompson5910 (jim_thompson5910):

nope, we got them all and we can't rewrite any more into the form y = kx so let's move onto inverse variation

jimthompson5910 (jim_thompson5910):

inverse variation: as one variable goes up, the other goes down (or vice versa)

jimthompson5910 (jim_thompson5910):

this can be represented in the general form y = k/x

jimthompson5910 (jim_thompson5910):

are there any equations that look like this

OpenStudy (anonymous):

Number 8?

OpenStudy (anonymous):

or number 4?

jimthompson5910 (jim_thompson5910):

close, but now that I look at it more closely, you can rewrite #8 as 36 = x/y 36y = x y = x/36 y = 1/36 times x So #8 is a direct variation

jimthompson5910 (jim_thompson5910):

# 4 is not an inverse variation because we're not dividing by any variables

OpenStudy (anonymous):

Eh. This stuff is kind of confusing.

jimthompson5910 (jim_thompson5910):

#6 is something that experiences inverse variation because notice how we're dividing by 't'

OpenStudy (anonymous):

Thats the only one isnt it?

jimthompson5910 (jim_thompson5910):

but there's more to the story as #6 also experiences direct variation (since there's a variable in the numerator)

jimthompson5910 (jim_thompson5910):

because #6 has both direct and inverse variation, this means that #6 is representing Combined Variation

OpenStudy (anonymous):

Which one would you pick?

jimthompson5910 (jim_thompson5910):

I would pick Combined Variation because that's a better description (it tells us more about the equation)

jimthompson5910 (jim_thompson5910):

so that would make #8 have Combined Variation as well (wasn't thinking about this one til just now)

jimthompson5910 (jim_thompson5910):

now onto joint variation, this is variation that's direct but it can have 2 or more variables on the right side So something like #4. y=6xz represents joint variation

OpenStudy (anonymous):

thats the only one isnt it?

jimthompson5910 (jim_thompson5910):

yes, it's the only one

jimthompson5910 (jim_thompson5910):

Sorry I'm a bit all over the place with this solution. To clean things up, the answers are.... 1. y=3x ....... direct 2. xy=14 ....... inverse (since we can rewrite into y = 14/x) 3. x^2=9 ....... no variation 4. y=6xz ....... joint 5. a=b/5 ...... direct 6. r=2s/t ....... combined 7. x-1=10 ..... no variation 8. 36=x/y ..... combined

jimthompson5910 (jim_thompson5910):

let me know if this helps and feel free to ask about anything in particular

OpenStudy (anonymous):

Thank you for helping. I have some more questions. If you can help me.

jimthompson5910 (jim_thompson5910):

sure, post to the left to make a new thread

jimthompson5910 (jim_thompson5910):

or if it's about these problems, then you can ask here

OpenStudy (anonymous):

Ok.

jimthompson5910 (jim_thompson5910):

let me know which option you're choosing

OpenStudy (anonymous):

new thread

jimthompson5910 (jim_thompson5910):

alright thanks